### v4

# You Only Sample Once: Taming One-Step Text-to-Image Synthesis by Self-Cooperative Diffusion GANs

Abstract
We introduce YOSO, a novel generative model designed for rapid, scalable, and high-fidelity one-step image synthesis. YOSO integrates the diffusion process with GANs to achieve the best of two worlds. Specifically, we smooth the distribution by the denoising generator itself, performing self-cooperative learning. We show that our method can serve as a one-step generation model training from scratch with competitive performance. Moreover, we show that our method can be extended to finetune pre-trained text-to-image diffusion for high-quality one-step text-to-image synthesis even with LoRA fine-tuning. In particular, we provide the first diffusion transformer that can generate images in one step trained on 512 resolution, with the capability of adapting to 1024 resolution without extra explicit training.
1 Introduction
Diffusion models (DMs) [46, 12, 50] have recently emerged as a powerful class of generative models, demonstrating state-of-the-art results in many generative modeling tasks, such as text-to-image [40, 60, 6, 7], text-to-video [4, 13], image editing [9, 5, 32] and controlled generation [64, 33]. However, the generation process of DMs requires iterative denoising, leading to slow generation speed. The sluggish generation time, coupled with the intensive computational requirements of large-scale DMs, constitutes a substantial barrier to their practical application and wider adoption.
Sampling from DMs can be regarded as solving a probability flow ordinary differential equation (PF-ODE) [50]. Some previous works [47, 27, 28, 2] focus on developing advanced ODE-solver, for reducing the sampling steps. However, they still require 20+ steps to achieve high-quality generation. Another line is distilling from pre-trained PF-ODEs [49, 26, 29, 30, 41], aiming to predict multi-step solution of PF-ODE solver by one step. Existing methods [29, 30] can achieve reasonable sample quality with 4+ steps. However, it is still challenging to generate high-quality samples with one step.
In contrast, generative adversarial networks (GANs) [8, 38] are naturally built on one-step generation with fast sampling speed. However, it is hard to extend GANs on large-scale datasets due to the serious training challenges, with worse sample quality compared to DMs [42, 17]. In this work, we present a novel approach to combine diffusion process and GANs. Previous works [54, 59, 60] have developed some variants of diffusion-GAN hybrid models. However, they rely on adding noise to smooth the adversarial divergence, yielding indirect learning in the one-step generation. To directly learn the one-step generation by adversarial divergence and to avoid unstable learning, we propose defining the one-step generation distribution on different levels of corrupted samples by diffusion process, where denotes the distribution of noisy data and denotes the denoised distribution parameterized by denoising generator. Then, we regard the one-step generation based on less corrupted samples as ground truth to perform adversarial divergence. This approach can be regarded as a self-cooperative process, as the generator learns from itself. Such an innovative design enables stable training and effective learning for one-step generation. Hence, we name our model YOSO representing You Only Sample Once. Moreover, we extend our YOSO to one-step text-to-image generation based on pre-trained models by several effective training techniques (i.e., latent perceptual loss and latent discriminator for efficient training along with the latent DMs; the informative prior initialization (IPI), and the quick adaption stage for fixing the noise scheduler).
Thanks to our effective designs, we can efficiently and effectively fine-tune existing pre-trained text-to-image DMs (i.e., stable diffusion [40] and PixArt- [6]) for high-quality one-step generation (See Fig. 1). Moreover, we are the first to support Low Rank Adaptation (LoRA) [16] fine-tuning for one-step generation of text-to-image.
Our work presents several significant contributions:
We introduce YOSO, a novel generative model that can generate high-quality images with one-step inference.
We propose several principled and effective training techniques enabling the fine-tuning of pre-trained text-to-image DMs for one-step text-to-image, requiring only ~10 A800 days.
We develop the first one-step diffusion transformer [36] with PixArt-, and the first LoRA enabling one-step generation, both of which will be released soon.
We conduct extensive experiments to demonstrate the effectiveness of the proposed YOSO, including image generation from scratch, text-to-image generation fine-tuning, compatibility with existing image customization, and image controllable modules.
2 Related Works
Recently, several works have been developed for few-step text-to-image generation. LCM [29] adapts consistency distillation [49] for stable diffusion (SD) and enables image generation in 4 steps with acceptable quality. InstaFlow [26] adapts rectified flows [25, 24] for SD, facilitating the generation of text-to-image in a mere single step. Despite this, the fidelity of the generated images is still poor. DMD [62] adapts variation score distillation [52] for SD, which requires training an additional DMs for the generator distribution, akin to the discriminator in GANs. Recently, combining DMs and GANs for one-step text-to-image generation has been explored. UFOGen [60] extends DDGANs [54] and SSIDMs [59] to SD by modifying the computation of reconstruction loss from corrupted samples to clean samples. However, it still performs adversarial matching using corrupted samples. ADD [44] introduces a one-step text-to-image generation based on SD. It follows earlier research [43] by employing a pre-trained image encoder, DINOv2 [34], as the backbone of the discriminator to accelerate the training. However, the discriminator design moves the training from latent space to pixel space, which substantially heightens computational demands. Moreover, they directly perform adversarial matches at clean real data, increasing the challenge of training. This requires the need for a better but expensive discriminator design and expensive R1 regularization to stabilize the training. In contrast, we perform the adversarial training by replacing real data with self-generated data to smooth the ground truth. Moreover, compared with UFOGen and ADD, our approaches can be trained from scratch to perform one-step generation, which is not demonstrated by them. Additionally, we extend our method not only to SD but also to PixArt- [6] which is based on diffusion transformer [36]. This demonstrates the wide application of our proposed YOSO.
3 Background
Diffusion Models. Diffusion models (DMs) [46, 12] define a forward process that gradually transforms samples from data distribution to Gaussian distribution by adding noise in steps with variance schedule : . The corrupted samples can be directly obtained by , where and . The parameterized reversed diffusion process is defined to gradually denoise: . The model can be trained by minimizing the negative ELBO [12, 20]: where is Gaussian posterior derived in [12]. A key assumption in diffusion is that the denoising step size from to is sufficiently small. This assumption ensures the true follows Gaussian distribution, enabling the effectiveness of modeling with Gaussian distribution.
Diffusion-GAN Hybrids. An issue in DMs is that the true does not follow Gaussian distribution when the denoising step size is not sufficiently small. Therefore, in order to enable large denoising step size, Diffusion GANs [54] propose to minimize the adversarial divergence between model and : , where and is imposed by a GAN generator. The capability of a GAN-based formulation, enables much larger denoising step sizes (i.e., 4 steps).
4 Method: Self-Cooperative Diffusion GANs
4.1 Revisiting of Diffusion GANs
A key issue in Diffusion-GAN hybrid models is that they match the generator distribution at corrupted data distribution. This formulation only indirectly learns the and , which are the distributions used for one-step generation, making the learning process less effective.
4.2 Our Design
To enable more effective learning for one-step generation, we propose to directly construct the learning objectives over clean data. We first construct a sequence distribution of clean data as follows:
where is the data distribution, and is the denoising generator. Note that is our denoising generator that predicts clean samples. If the network is parameterized to predict noise, we have .
Given the constructed distribution, we can formulate the optimization objective as follows:
where and . The optimization objective is constructed by the combination of adversarial divergence and KL divergence. Specifically, the adversarial divergence focuses on matching over the distribution level, ensuring the generation quality, and the KL divergence focuses on matching over point level, ensuring the mode cover.
However, it is hard to directly learn the adversarial divergence over clean data distribution, akin to the challenges with GAN training. To tackle these challenges, previous Diffusion GANs [54, 59] have pivoted towards learning the adversarial divergence over the corrupted data distribution. Unfortunately, as analyzed before, such an approach fails to directly match , curtailing the efficacy of one-step generation. Moreover, it also compels the discriminator to fit different levels of noise, leading to limited capability.
Recall that is defined as . The quality of the distribution has two key factors: 1) the ability of the trainable generator ; 2) the information given by .
Hence given the generator fixed, if we increase the information in , it is supposed that we can get a better distribution. In other words, it is highly likely that is superior to . Motivated by the cooperative approach [55, 56, 57, 11] which uses an MCMC revised version of model distribution to learn the generator, we suggest to use as ground truth to learn , which constructs the following training objective:
where denotes stop-gradient operator. This training objective can be regarded as a self-cooperative approach, since the ‘revised’ samples from and samples from are generated by the same network . Note that we only replace data distribution as in the adversarial divergence for smoothing the learning objective, as recent work [31] has found that it is beneficial to learn the generator with a mix of real data and revised data.
In the distribution matching objective above, we apply the non-saturating GAN objective to minimize the adversarial divergence of marginal distributions. And the KL divergence for point matching can be optimized by loss. Hence a tractable training objective is formulated as follows:
where is the discriminator network. We find that the self-cooperative approach is connected with Consistency Training [49]. However, Consistency Training considers the as an approximated ODE solution of to perform a point-to-point match. In contrast, our proposed objective matches and in marginal distribution level, which avoids the approximated error of ODE.
To further ensure the mode cover of the proposed model, we can add consistency loss to our objective as regularization, which constructs our final training objective:
5 Towards One-Step Text-to-Image Synthesis
Since training a text-to-image model from scratch is quite expensive, we suggest using pre-trained text-to-image DMs as initialization with Self-Coopeartive Diffusion GANs. In this section, we introduce several principled designs for developing a generation model that enables one-step text-to-image synthesis based on pre-trained DMs.
5.1 Using pre-trained models for Training
Latent Perceptual Loss. Prior research [15, 14, 49] has confirmed the effectiveness of perceptual loss in various domains. Notably, recent studies [26, 49] has found that the LPIPS loss [65] is crucial for obtaining few-step DMs with high sample quality. However, a notable drawback is that the LPIPS loss is computed in the data space, which is expensive. In contrast, the popular SD operate in the latent space to reduce computational demands. Hence, using LPIPS loss in training latent DMs is pretty expensive, as it requires not only the computation of LPIPS loss in the data space but also an additional decoding operation. Realizing that the pre-trained SD can serve as an effective feature extractor [58], we suggest using pre-trained SD to perform latent perceptual loss. However, SD is a UNet, whose final layer predicts epsilon with dimensions identical to the data. Hence we propose using the bottleneck layer of the UNet for computing:
where is the latent encoded images by VAE and is the text feature. We note that the benefit of computing latent perceptual loss by SD is not only the computational efficiency but also the incorporation of text features, which are crucial in text-to-image tasks.
Latent Discriminator. Training GANs for the text-to-image on large-scale datasets faces serious challenges. Specifically, unlike unconditional generation, the discriminator for the text-to-image task should justify based on both image quality and text alignment. This challenge is more obvious during the initial stage of training. To address this issue, previous pure GANs [18] for text-to-image propose complex learning objectives and require expensive costs for training. The learning of GANs has been shown to benefit from using a pre-trained network as the discriminator. As discussed above, the pre-trained SD has learned representative features. Hence, we suggest applying the pre-trained SD for constructing the latent discriminator. Similar to latent perceptual loss, we only use half UNet for the discriminator followed by a simple predict head. The advantages of the proposed strategy are twofold: 1) we use the informative pre-trained network as initialization; 2) the discriminator is defined over latent space, which is computationally efficient. Unlike previous work [44] that defines a discriminator over data space which required decode latent and backward from decoder, yielding expensive computational cost. By applying the Latent Discriminator, we observe a stable training process with fast convergence. Note that, expensive R1 regularization is not needed in our approach.
Decoupled Scheduler. We find that the optimal scheduler for performing consistency loss and adversarial loss are not identical. Hence, to better unleash the capability of pre-trained DMs, we propose using decoupled schedulers to construct the following loss in fine-tuning pre-trained DMs:
where , , and we let and in experiments.
5.2 Fixing the Noise Scheduler
A common issue in DMs is that the final corrupted samples are not pure noise. For example, the noise scheduler used by the SD makes the corrupted samples at the final timestep as: , where the terminal Signal-to-noise ratio (SNR) is , effectively creates a gap between training and inference. Previous work [23] has only observed that this makes DMs unable to produce pure black or white images. However, we find that this issue yields serious problems in one-step generation. As shown in Fig. 2, there are notable artifacts in one-step generation if we directly sample noise from standard Gaussian distribution. The reason may be that in multi-step generation, the gap can be gradually fixed in sampling, while one-step generation reflects the gap more directly. To address this issue, we provide two simple yet effective solutions.
Informative Prior Initialization (IPI). The non-zero terminal SNR issue is similar to prior hole issues in VAEs [22, 3, 21]. Therefore, we can use informative prior instead of non-informative prior to effectively address this issue. For simplicity, we adopt a learnable Gaussian Distribution , whose optimal formulation is given below: , where and can be efficiently estimated by finite samples, and follows standard Gaussian distribution. As shown in Fig. 2, the artifacts in one-step generation are immediately removed after applying IPI. We note that the performance is achieved with minimal adjustment, enabling the possibility of developing one-step text-to-image generation by LoRA fine-tuning.
Quick Adaption to V-prediction and Zero Terminal SNR. The IPI suffers from numerical instability when terminal SNR is pretty low. As shown in Fig. 3, we fine-tune PixArt- [6] whose terminal SNR is 4e-5 with introduced technique and -prediction, failing in one-step generation. Following [23], we suggest switching to v-prediction [41] and zero terminal SNR. However, we found that direct transition convergences slowly (see Appendix F). This is unacceptable in solving large-scale text-to-image with limited computational resources. To address this, we propose a quick adapt-stage:
Adapt-stage-I switch to v-prediction : where , , denotes the desired v-prediction model, and denotes the frozen pre-trained model.
Adapt-stage-II switch to zero terminal SNR : Note that in this stage, we only change the scheduler to zero terminal SNR for student model. This avoids numerical instability issues of -prediction with zero terminal SNR. We note that the zero terminal SNR scheduler has lower SNR than the original schedule at each timestep, formulating an effective distillation objective.
We note that this adapt-stage converges rapidly, typically requiring only 1k iterations to initialize training YOSO. This enables a quick adaption to v-prediction and zero terminal SNR from pre-trained -prediction DMs, thereby comprehensively mitigating the non-zero terminal SNR issue in the noise scheduler.
6 Experiments
In this section, we evaluate the performance of the proposed YOSO. In Sec. 6.1, we evaluate YOSO on unconditional image generation by training from scratch, and we conduct ablation studies in Sec. 6.2 to highlight the effectiveness of our algorithm. Sec. 6.3 examines YOSO in the context of text-to-image generation by fine-tuning pre-trained PixArt- [6] and Stable Diffusion [40]. In Sec. 6.4, we show that the YOSO can be used for several downstream applications.
Model | FID | NFE | Time (s) |
YOSO | 3.82 | 1 | 0.05 |
Ablation: | |||
YOSO w/o consistency loss | 4.41 | 1 | 0.05 |
YOSO w/o LPIPS loss | 4.63 | 1 | 0.05 |
YOSO w/ | 10.85 | 1 | 0.05 |
UFOGen [60] | 45.15 | 1 | 0.05 |
DDGANs [54] | 3.75 | 4 | 0.21 |
CT [49] | 8.70 | 1 | - |
iCT [48] | 2.83 | 1 | - |
DDPM [12] | 3.21 | 1000 | 80.5 |
EDM [19] | 2.04 | 36 | - |
DDIM [47] | 4.67 | 50 | 4.01 |
StyleGAN2 w/o ADA | 8.32 | 1 | 0.04 |
StyleGAN2 w/ADA | 2.92 | 1 | 0.04 |
6.1 Unconditional Image Generation
Experiment Setting. We use the same UNet and discriminator architecture as used in DDGANs [54]. We set the denoising steps as four. For the evaluation metric, we choose the Fréchet inception distance (FID) [10]. We train the proposed model on the CIFAR-10 dataset [63]. Following Consistency Models [49], we employ LPIPS loss instead of MSE to compute image distance.
Results. The quantitative results are shown in Tab. 1. We observe that our method provides competitive performance compared with the best DMs and GANs. Although some variants of DMs achieve better results, they generally require much longer sample time compared to ours. We also find that among variants of GANs, our performance is surpassed only marginally by StyleGAN2 employing adaptive data augmentation. We also observe that the consistency models provide competitive and even better results compared to us. However, we note that the training of consistency models requires tons of computational resources and careful design of the noise schedule. As shown in Fig. 4, the consistency models require much more computational resources than ours, including more parameters, larger batch size, and more training iterations. To ensure a fair comparison, we re-produce the consistency models by using the same batch size and training iterations as ours. Based on the fair setting, we find that the performance of consistency models significantly decrease, showing way worse performance compared to YOSO. We highlight that the low demand for computational resources enables us to scale the proposed YOSO to large-scale text-to-image datasets.
6.2 Ablation Studies
We provide additional insight into the effectiveness of the proposed YOSO by performing ablation studies to some potential variants on CIFAR-10 dataset.
Effect of Consistency Loss and LPIPS Loss. We evaluate the effect of consistency loss and LPIPS loss in YOSO. We find that both consistency loss and LPIPS loss can improve the image quality, as indicated by FID (see Tab. 1). However, we notice that if removing consistency loss or LPIPS loss, the generation performance would not decrease significantly. This indicates the effectiveness of our proposed adversarial formulation.
Adversarial Divergence . The key design of our proposed YOSO is smoothing the distribution by using self-generated data instead of real data as ground truth to perform adversarial divergence. We evaluate the variant of using as adversarial divergence. As shown in Tab. 1, the performance of the variant is significantly worse than YOSO. Moreover, besides sample quality, we find that the training of our proposed YOSO is much more stable than the variant. As shown in Fig. 4, the discriminator loss of the variant is not stable and is much smaller than ours. This indicates that our approach can effectively smooth the distribution to make the distributions of fake and real samples closer to each other, making discrimination harder and hence resulting in higher discriminator loss. Furthermore, we observe mode collapse in the training of the variant. We note that a similar case also happened in SDXL-Turbo, a strong one-step text-to-image model, which also performs adversarial divergence directly against real data (See Fig. 5). These results indicate the superiority of the proposed YOSO.
Comparison to UFOGen. Similar to us, UFOGen does not directly perform adversarial divergence against real data either. However, they perform adversarial divergence over corrupted data (i.e., ). We argue that the formulation is less effective in constructing one-step generation model since their adversarial divergence only indirectly matches the one-step generation at clean data. We implement the UFOGen on CIFAR-10 with our own. Note that we also perform consistency loss in the UFOGen variant to ensure a fair comparison, as the consistency loss has been found useful in YOSO. As shown in Tab. 1, UFOGen only obtains 45.15 FID on CIFAR-10, significantly worse than YOSO with 3.82 FID. This further demonstrates the superiority of our proposed YOSO. Through comparison with different adversarial divergence formulations, we find that our formulated adversarial divergence not only smooths the distribution to stabilize the training but also provides effective learning on one-step generation.
6.3 Text-to-Image Generation
Experiment Setting. We initialize the GAN generator by pre-trained PixArt- which is a diffusion transformer with 0.6B parameters. We note that 0.6B parameters can be demonstrated as lightweight among text-to-image DMs. For the GAN discriminator, we construct the latent discriminator by pre-trained SD 1.5 using the proposed construction introduced in Sec. 5.1. We switch the pre-trained PixArt- to v-prediction by the proposed technique introduced in Sec. 5.1, followed by training on the JourneyDB dataset [35] with resizing as 512 resolution. We apply a batch-size of 256 and a constant learning rate of 2e-5 during training. We observe the training convergence fast, requiring only 30k iterations and around 10 A800 days to be trained.
Evaluation. For image quality, we employ Aesthetic Score (AeS) [45] to evaluate. We also adopt the Human Preference Score (HPS) v2.1 [53] to evaluate the image-text alignment and human preference. The AeS is computed by an aesthetic score predictor trained on LAION [45] datasets, without considering the image-text alignment. HPS is trained to predict human preference given image-text pairs, considering the image-text alignment and human aesthetic. We mainly compare our model against the open-source state-of-the-art (SOTA) models, i.e., SD-Turbo [44] and SDXL-Turbo [44].
Quantitative Results. The quantitative results are presented in Tab. 3. To provide a direct comparison with SD-Turbo, we also fine-tune a model based on pre-trained SD 1.5 on JourneyDB with the introduced technique. We would like to highlight that it is better to conduct training on LAION dataset [45] which is the training set used by SD 1.5. However, it is not available when we conducted this research. Fine-tuning SD 1.5 on other datasets suffers from distribution shift, facing a risk of catastrophic forgetting. As shown in Tab. 3, both YOSO-SD and YOSO-PixArt- clearly beat the SD-Turbo which is based on SD 2.1, a better SD compared to SD 1.5. It is also observed that YOSO-PixArt- outperforms SDXL-Turbo which is a much larger model with 2.6B parameters.
Qualitative Comparison. To provide more comprehensive insight in understanding the performance of our YOSO, we further provide the qualitative comparison in Fig. 5. The comparison shows that: YOSO clearly beats SD-Turbo in terms of both image quality and prompt alignment. Specifically, the samples by SD-Turbo are blurry to some extent. Although SDXL-Turbo with a larger model size improves the sample quality compared to SD-Turbo, we find that the samples by YOSO are sharper than SDXL-Turbo to some extent, and better in prompt alignment. For example, in the first row of Fig. 5 related to a Pikachu, both SD-Turbo and SDXL-Turbo cannot follow the prompt to generate angry expressions and red eyes, while YOSO follows the prompt more precisely. Moreover, it is important to highlight that SDXL-Turbo seems to have a serious issue in mode collapse. Specifically, the third row of Fig. 5 shows that samples by SDXL-Turbo are really close to each other. Overall, based on quantitative and qualitative results, we conclude that our YOSO is better in sample quality, prompt alignment, and mode cover compared with the SOTA one-step text-to-image models.
6.3.1 Zero-Shot One-Step 1024 Resolution Generation
Due to the computational resource limitation, we trained YOSO on 512 resolution. However, there is a 1024-resolution version of PixArt-, obtained by continuing training on the 512-resolution version of PixArt-. Motivated by the effectiveness of the LoRA combination [30], we suggest constructing a similar combination as follows:
where and means the model weight. We let in our experiments. As shown in Fig. 1(b), the YOSO constructed in this way can generate high-quality images with 1024 resolution. Note that YOSO even changes the predicted objective to v-prediction. The impressive performance indicates the robust generalization ability of our proposed YOSO.
6.3.2 Fine-tuning Stable Diffusion via LoRA
We further conduct experiments on fine-tuning SD 1.5 via LoRA to show the effectiveness of proposed YOSO. As we perform LoRA fine-tuning, it is expected that it will be worse than full fine-tuning, we propose to use SDXL-Turbo to generate data for LoRA fine-tuning, which avoids the risk of distribution shift to some extent and is economically efficient. We call the model fine-tuned in this way as YOSO-LoRA. We only generate one image for each caption, avoiding the risk of mode collapse brought by SDXL-Turbo. We train YOSO-LoRA with only 64 batch size and 4000 iterations. We present the quantitative results in Tab. 4 and qualitative results in Fig. 6. It can be observed that YOSO-LoRA achieves even better results than LCM with 4 steps and is comparable to InstaFLow with full fine-tuning. And from Fig. 6, YOSO-LoRA has a clearly better sample quality than LCM with 2 steps and InstaFlow. These results indicate that our algorithm enables one-step text-to-image synthesis with reasonable quality by LoRA fine-tuning.
6.4 Application
One promising and attractive property of DMs is that they can be used in multiple downstream tasks. In this section, we show the capability of our proposed YOSO in various downstream tasks, keeping the unique advantages of DMs. Since the relevant applications are mostly developed for SD, we mainly show the results obtained by YOSO-LoRA.
Image-to-image. As shown in Fig. 6, our YOSO-LoRA is capable of performing high-quality image-to-image editing [32] in one step.
Compatibility with ControlNet and Different Base Models. As shown in Fig. 6, our YOSO-LoRA is compatible with ControlNet [64], following the condition well. Our YOSO-LoRA is also compatible with different base models (e.g., dreamshaper and realistic vision) fine-tuned from SD 1.5, preserving their style well. We find that when applying YOSO-LoRA to new base models, it fails to produce reasonable samples in one step. This may be due to the capability of LoRA and the distribution shift of training data.
Model | NFE | HPS | AeS |
SD 2.1-Turbo [44] | 1 | 27.06 | 5.83 |
Realistic-vision | 50 | 29.76 | 6.35 |
w/ YOSO-LoRA | 4 | 29.50 | 6.21 |
w/ YOSO-LoRA-v2 | 1 | 27.13 | 6.12 |
w/ YOSO-LoRA-v2 | 2 | 29.41 | 6.33 |
Dreamshaper | 1 | 26.64 | 6.08 |
w/ YOSO-LoRA-v2 |
Enabling Data Flywheel. YOSO-LoRA trained on the data generated by SDXL-Turbo, fails to generate reasonable samples in one step when compatible with new base models. In light of the fact that combining YOSO-LoRA with new base models can generate high-quality samples in 4 steps with minimal performance loss as well as preserved style, we propose using YOSO-generated data for training YOSO-LoRA based on existing community models (e.g., Realistic-vision). We call the trained model YOSO-LoRA-v2. As shown in Tab. 5, YOSO-LoRA-v2 can even outperform the full fine-tuned SD-2.1-Turbo in one step. This shows that the YOSO can be trained on slightly shifted data first, and then applied to generate high-quality un-shifted data to train a better YOSO. Moreover, the YOSO-LoRA-v2 can be compatible with some unseen base models (e.g., Dreamshaper) for one-step generation (see Figs. 6 and 5). This shows that combining YOSO-LoRA with community models can generate unshifted training data, enabling the YOSO-LoRA-v2 to focus on learning the acceleration with better compatibility.
7 Conclusion
In summary, we present YOSO, a new generative model enabling high-quality one-step generation. Our novel designs combine the diffusion process and GANs, enabling not only one-step generation training from scratch but also one-step text-to-image generation fine-tuning from pre-trained models. We will release our model to advance the research of text-to-image synthesis.
Appendix A Broader Impacts
This work presents YOSO, a method that accelerates multi-step large-scale text-to-image diffusion models into a one-step generator. On the positive side, while this is academic research, we believe the proposed YOSO can be widely applied in industry, and its high efficiency can lead to energy savings and environmental benefits. However, when these rapid generation models are manipulated by malicious actors, they can also simplify and accelerate the creation of harmful information. Although our work focuses on scientific research, we will take actions to reduce the harmful information, such as filtering out harmful content in the dataset.
Appendix B Limitation
Our model, like most text-to-image diffusion models, may exhibit shortcomings in terms of fairness, as well as in handling specific details and accurately controlling the number of targets. We plan to explore these unresolved issues in the generation field in our future work, in order to enhance the model’s capabilities in text generation, fairness, detail control, and quantitative control.
Appendix C Safeguards
The YOSO-SD and YOSO-PixArt- are trained on the JourneyDB dataset [35], which has undergone rigorous human and machine-based filtering to ensure that there are no harmful or violent images in the dataset. YOSO-LoRA and YOSO-LoRA-V2, on the other hand, were trained using data generated by SDXL-turbo and YOSO-LoRA, respectively. To reduce the risk associated with the models, we utilized the built-in safety checker from diffusers [51] to filter out harmful data.
Appendix D Training Algorithm
We present the algorithm for training YOSO from scratch in Algorithm 1.
Appendix E Additional Related Work
Text-to-image Diffusion Models. Since Diffusion models have shown a stable training process and are well-suited for scaling up generative models, numerous works have been proposed to extend DMs for text-to-image generation [39, 1, 40, 61]. Latent DMs [40, 37] are widely adopted for high-resolution image and video generation due to their computational efficiency.
Appendix F The Effect of Quick-Adaption
We present a qualitative comparison between YOSO w/o quick adaption and YOSO w/ quick adaption in Fig. 7. Note that the YOSO and variants here are trained with only 5k iterations. And for YOSO w/ quick adaption, we incorporate the 1k iterations used for quick adaptation into the total 5k training iterations to ensure a fair comparison. As shown in Fig. 7, YOSO w/ quick adaption demonstrates significantly better image quality compared to YOSO w/o quick adaption. The generated images of YOSO w/o quick adaption exhibit over-smoothing and severe artifacts. This indicates that the convergence of direct transition to v-prediction and zero terminal SNR is relatively slower. Additionally, it is worth highlighting that although there are only 5k training iterations, the one-step generation quality of YOSO is already reasonable, which demonstrates the effectiveness and quick convergence of the proposed quick adaption stage and YOSO.
Appendix G Experiment Setting Details
G.1 Unconditional Generation Experiments
For the generator, we use the Adam optimizer with = 0.9 and = 0.999; for the discriminator, we use the Adam optimizer with = 0. and = 0.999. We adopt a constant learning rate of 2e-4 for both the discriminator and the generator. We apply EMA with a coefficient of 0.9999 for the generator. We let the and .
G.2 Text-To-Image Experiments
For the generator, we use the AdamW optimizer with = 0.9 and = 0.999; for the discriminator, we use the AdamW optimizer with = 0. and = 0.999. We adopt a constant learning rate of 2e-5 for both the discriminator and the generator. We apply gradient norm clipping with a value of 1.0 for the generator only. We use batch size 256. For full fine-tuning, we apply EMA with a coefficient of 0.9999 for the generator. For LoRA fine-tuning, we apply EMA with a coefficient of 0.999 for the generator. Generally, the training is done within 30k iterations for full fine-tuning, while within 5k iterations for LoRA fine-tuning. We let the and .

### v5

@@
 Abstract
-We introduce YOSO, a novel generative model designed for rapid, scalable, and high-fidelity one-step image synthesis.
-YOSO integrates the diffusion process with GANs to achieve the best of two worlds.
-Specifically, we smooth the distribution by the denoising generator itself, performing self-cooperative learning.
+Recently, some works have tried to combine diffusion and Generative Adversarial Networks (GANs) to alleviate the computational cost of the iterative denoising inference in Diffusion Models (DMs).
+However, existing works in this line suffer from either training instability and mode collapse or subpar one-step generation learning efficiency.
+To address these issues, we introduce YOSO, a novel generative model designed for rapid, scalable, and high-fidelity one-step image synthesis with high training stability and mode coverage.
+Specifically, we smooth the adversarial divergence by the denoising generator itself, performing self-cooperative learning.
 We show that our method can serve as a one-step generation model training from scratch with competitive performance.
-Moreover, we show that our method can be extended to finetune pre-trained text-to-image diffusion for high-quality one-step text-to-image synthesis even with LoRA fine-tuning.
-In particular, we provide the first diffusion transformer that can generate images in one step trained on 512 resolution, with the capability of adapting to 1024 resolution without extra explicit training.
+Moreover, we extend our YOSO to one-step text-to-image generation based on pre-trained models by several effective training techniques (i.e., latent perceptual loss and latent discriminator for efficient training along with the latent DMs; the informative prior initialization (IPI), and the quick adaption stage for fixing the flawed noise scheduler).
+Experimental results show that YOSO achieves the state-of-the-art one-step generation performance even with Low-Rank Adaptation (LoRA) fine-tuning.
+In particular, we show that the YOSO-PixArt- can generate images in one step trained on 512 resolution, with the capability of adapting to 1024 resolution without extra explicit training, requiring only ~10 A800 days for fine-tuning.
+Our code is provided at https://github.com/Luo-Yihong/YOSO.
 1 Introduction
-Diffusion models (DMs) [46, 12, 50] have recently emerged as a powerful class of generative models, demonstrating state-of-the-art results in many generative modeling tasks, such as text-to-image [40, 60, 6, 7], text-to-video [4, 13], image editing [9, 5, 32] and controlled generation [64, 33].
+Diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) have recently emerged as a powerful class of generative models, demonstrating state-of-the-art results in many generative modeling tasks, such as text-to-image (Rombach et al., 2022; Xu et al., 2023c; Chen et al., 2024; Feng et al., 2023), text-to-video (Blattmann et al., 2023; Hong et al., 2022), image editing (Hertz et al., 2022; Brooks et al., 2023; Meng et al., 2022) and controlled generation (Zhang et al., 2023; Mou et al., 2023).
 However, the generation process of DMs requires iterative denoising, leading to slow generation speed.
-The sluggish generation time, coupled with the intensive computational requirements of large-scale DMs, constitutes a substantial barrier to their practical application and wider adoption.
-Sampling from DMs can be regarded as solving a probability flow ordinary differential equation (PF-ODE) [50].
-Some previous works [47, 27, 28, 2] focus on developing advanced ODE-solver, for reducing the sampling steps.
+Coupled with the intensive computational requirements of large-scale DMs, they constitute a substantial barrier to their practical application and wider adoption.
+Sampling from DMs can be regarded as solving a probability flow ordinary differential equation (PF-ODE) (Song et al., 2021).
+Some previous works (Song et al., 2020; Lu et al., 2022a; b; Bao et al., 2022) focus on developing advanced ODE-solvers, for reducing the sampling steps.
 However, they still require 20+ steps to achieve high-quality generation.
-Another line is distilling from pre-trained PF-ODEs [49, 26, 29, 30, 41], aiming to predict multi-step solution of PF-ODE solver by one step.
-Existing methods [29, 30] can achieve reasonable sample quality with 4+ steps.
+Another line is distilling from pre-trained PF-ODEs (Song et al., 2023; Liu et al., 2023; Luo et al., 2023a; b; c; Salimans & Ho, 2022), aiming to predict multi-step solution of PF-ODE solver by one step.
+Existing methods (Luo et al., 2023a; b) can achieve reasonable sample quality with 4+ steps.
 However, it is still challenging to generate high-quality samples with one step.
-In contrast, generative adversarial networks (GANs) [8, 38] are naturally built on one-step generation with fast sampling speed.
-However, it is hard to extend GANs on large-scale datasets due to the serious training challenges, with worse sample quality compared to DMs [42, 17].
+In contrast, generative adversarial networks (GANs) (Goodfellow et al., 2014; Radford et al., 2016) are naturally built on one-step generation with fast sampling speed.
+However, it is hard to extend GANs on large-scale datasets due to training challenges (Sauer et al., 2022; Kang et al., 2023), resulting in worse sample quality compared to DMs (Sauer et al., 2023b; Kang et al., 2023).
 In this work, we present a novel approach to combine diffusion process and GANs.
-Previous works [54, 59, 60] have developed some variants of diffusion-GAN hybrid models.
-However, they rely on adding noise to smooth the adversarial divergence, yielding indirect learning in the one-step generation.
-To directly learn the one-step generation by adversarial divergence and to avoid unstable learning, we propose defining the one-step generation distribution on different levels of corrupted samples by diffusion process, where denotes the distribution of noisy data and denotes the denoised distribution parameterized by denoising generator.
-Then, we regard the one-step generation based on less corrupted samples as ground truth to perform adversarial divergence.
-This approach can be regarded as a self-cooperative process, as the generator learns from itself.
+The key to achieving access is finding a way to smooth the adversarial divergence for stabilize the training while maintaining effective one-step learning.
+Previous works (Xiao et al., 2022; Xu et al., 2023b; c; Sauer et al., 2023c) have developed some variants for combining diffusion and GANs.
+However, existing works either directly perform adversarial divergence against real data without smoothing strategy which suffers from unstable training and mode collapse, or rely on adding noise to smooth the adversarial divergence to stabilize the training which delivers less effective one-step generation learning.
+To enjoy the best of both worlds, we propose to smooth the adversarial divergence by denoising the generator itself.
+In particular, we regard the one-step denoising generation based on less corrupted samples as ground truth and regard the one-step denoising generation based on more corrupted samples as student distribution to perform adversarial divergence.
+The approach can not only naturally reduce the distance between target distribution and student distribution to stabilize the training, but also form effective one-step learning on clean samples.
+The learning process can be viewed as a self-cooperative process (Xie et al., 2018), as the generator learns from itself.
 Such an innovative design enables stable training and effective learning for one-step generation.
-Hence, we name our model YOSO representing You Only Sample Once.
-Moreover, we extend our YOSO to one-step text-to-image generation based on pre-trained models by several effective training techniques (i.e., latent perceptual loss and latent discriminator for efficient training along with the latent DMs; the informative prior initialization (IPI), and the quick adaption stage for fixing the noise scheduler).
-Thanks to our effective designs, we can efficiently and effectively fine-tune existing pre-trained text-to-image DMs (i.e., stable diffusion [40] and PixArt- [6]) for high-quality one-step generation (See Fig.
+Hence, we name our model YOSO, short for You Only Sample Once.
+Moreover, we extend our YOSO to one-step text-to-image generation based on pre-trained models and introduce several effective training techniques (i.e., latent perceptual loss and latent discriminator for efficient training along with the latent DMs; the informative prior initialization (IPI), and the quick adaption stage for fixing the noise scheduler).
+Thanks to our effective designs, we can efficiently and effectively fine-tune existing pre-trained text-to-image DMs (i.e., stable diffusion (Rombach et al., 2022) and PixArt- (Chen et al., 2024)) for high-quality one-step generation (see Fig.
 1).
-Moreover, we are the first to support Low Rank Adaptation (LoRA) [16] fine-tuning for one-step generation of text-to-image.
+Furthermore, we are the pioneers in supporting Low-Rank Adaptation (LoRA) (Hu et al., 2022) fine-tuning in one-step text-to-image generation for enhanced efficiency, delivering state-of-the-art performance.
 Our work presents several significant contributions:
-We introduce YOSO, a novel generative model that can generate high-quality images with one-step inference.
-We propose several principled and effective training techniques enabling the fine-tuning of pre-trained text-to-image DMs for one-step text-to-image, requiring only ~10 A800 days.
-We develop the first one-step diffusion transformer [36] with PixArt-, and the first LoRA enabling one-step generation, both of which will be released soon.
-We conduct extensive experiments to demonstrate the effectiveness of the proposed YOSO, including image generation from scratch, text-to-image generation fine-tuning, compatibility with existing image customization, and image controllable modules.
-2 Related Works
-Recently, several works have been developed for few-step text-to-image generation.
-LCM [29] adapts consistency distillation [49] for stable diffusion (SD) and enables image generation in 4 steps with acceptable quality.
-InstaFlow [26] adapts rectified flows [25, 24] for SD, facilitating the generation of text-to-image in a mere single step.
-Despite this, the fidelity of the generated images is still poor.
-DMD [62] adapts variation score distillation [52] for SD, which requires training an additional DMs for the generator distribution, akin to the discriminator in GANs.
-Recently, combining DMs and GANs for one-step text-to-image generation has been explored.
-UFOGen [60] extends DDGANs [54] and SSIDMs [59] to SD by modifying the computation of reconstruction loss from corrupted samples to clean samples.
-However, it still performs adversarial matching using corrupted samples.
-ADD [44] introduces a one-step text-to-image generation based on SD.
-It follows earlier research [43] by employing a pre-trained image encoder, DINOv2 [34], as the backbone of the discriminator to accelerate the training.
-However, the discriminator design moves the training from latent space to pixel space, which substantially heightens computational demands.
-Moreover, they directly perform adversarial matches at clean real data, increasing the challenge of training.
-This requires the need for a better but expensive discriminator design and expensive R1 regularization to stabilize the training.
-In contrast, we perform the adversarial training by replacing real data with self-generated data to smooth the ground truth.
-Moreover, compared with UFOGen and ADD, our approaches can be trained from scratch to perform one-step generation, which is not demonstrated by them.
-Additionally, we extend our method not only to SD but also to PixArt- [6] which is based on diffusion transformer [36].
-This demonstrates the wide application of our proposed YOSO.
+We introduce YOSO, a novel generative model that can generate high-quality images with one-step inference, with a stable training process and good mode coverage.
+We further scale up YOSO by several principled and effective proposed training techniques that enable low-resource fine-tuning of pre-trained text-to-image DMs for one-step text-to-image, requiring only ~10 A800 days.
+We conduct extensive experiments to demonstrate the effectiveness of the proposed YOSO, including image generation from scratch, text-to-image generation fine-tuning, compatibility with existing image customization and image controllable modules.
 2 Background
 Diffusion models.
-Diffusion models (DMs) [46, 12] define a forward process that gradually transforms samples from data distribution to Gaussian distribution by adding noise in steps with variance schedule : .
+Diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020) define a forward process that gradually transforms samples from data distribution to Gaussian distribution by adding noise in steps with variance schedule : .
 The corrupted samples can be directly obtained by , where and .
 The parameterized reversed diffusion process is defined to gradually denoise: .
-The model can be trained by minimizing the negative ELBO [12, 20]: where is Gaussian posterior derived in [12].
+The model can be trained by minimizing the negative ELBO (Ho et al., 2020; Kingma et al., 2021): where is Gaussian posterior derived in (Ho et al., 2020).
 A key assumption in diffusion is that the denoising step size from to is sufficiently small.
@@
 An issue in DMs is that the true does not follow Gaussian distribution when the denoising step size is not sufficiently small.
-Therefore, in order to enable large denoising step size, Diffusion GANs [54] propose to minimize the adversarial divergence between model and : , where and is imposed by a GAN generator.
-The capability of a GAN-based formulation, enables much larger denoising step sizes (i.e., 4 steps).
+Therefore, in order to enable large denoising step size, Diffusion GANs (Xiao et al., 2022) propose to minimize the adversarial divergence between model and : , where and is imposed by a GAN generator.
+The capability of a GAN-based formulation enables much larger denoising step sizes (i.e., 4 steps).
 3 Method: Self-Cooperative Diffusion GANs
-4.1 Revisiting of Diffusion GANs
-A key issue in Diffusion-GAN hybrid models is that they match the generator distribution at corrupted data distribution.
+A key issue in Diffusion-GAN hybrid models (Xiao et al., 2022; Xu et al., 2023b; c) is that they match the generator distribution with the corrupted data distribution.
 This formulation only indirectly learns the and , which are the distributions used for one-step generation, making the learning process less effective.
@@
 where and .
-The optimization objective is constructed by the combination of adversarial divergence and KL divergence.
-Specifically, the adversarial divergence focuses on matching over the distribution level, ensuring the generation quality, and the KL divergence focuses on matching over point level, ensuring the mode cover.
+The optimization objective is constructed by combining adversarial divergence and KL divergence.
+Specifically, the adversarial divergence focuses on matching over the distribution level, ensuring the generation quality, and the KL divergence focuses on matching over point level, ensuring the mode coverage.
 However, it is hard to directly learn the adversarial divergence over clean data distribution, akin to the challenges with GAN training.
-To tackle these challenges, previous Diffusion GANs [54, 59] have pivoted towards learning the adversarial divergence over the corrupted data distribution.
+To tackle these challenges, previous Diffusion GANs (Xiao et al., 2022; Xu et al., 2023b) have pivoted towards learning the adversarial divergence over the corrupted data distribution.
 Unfortunately, as analyzed before, such an approach fails to directly match , curtailing the efficacy of one-step generation.
@@
 Hence given the generator fixed, if we increase the information in , it is supposed that we can get a better distribution.
-In other words, it is highly likely that is superior to .
-Motivated by the cooperative approach [55, 56, 57, 11] which uses an MCMC revised version of model distribution to learn the generator, we suggest to use as ground truth to learn , which constructs the following training objective:
-where denotes stop-gradient operator.
+In other words, it is highly likely that is superior to , where .
+Motivated by the cooperative approach (Xie et al., 2018; 2021; 2022; Hill et al., 2022) which uses an MCMC revised version of model distribution to learn the generator, we suggest to use as ground truth to learn , which constructs the following training objective:
+where denotes stop-gradient operator and the second term is named as cooperative adversarial loss.
 This training objective can be regarded as a self-cooperative approach, since the ‘revised’ samples from and samples from are generated by the same network .
-Note that we only replace data distribution as in the adversarial divergence for smoothing the learning objective, as recent work [31] has found that it is beneficial to learn the generator with a mix of real data and revised data.
+Note that we only replace data distribution as in the adversarial divergence for smoothing the learning objective, as recent work (Luo et al., 2023d) has found that it is beneficial to learn the generator with a mix of real data and revised data.
+We briefly verify the theoretical rationality of the cooperative adversarial divergence below.
+Proposition 1
+The optimal solution of the cooperative adversarial loss reaches .
+This proposition tells that the proposed cooperative adversarial loss can recover the real data distribution when the network’s capability is sufficiently large, demonstrating the theoretical rationality of the proposed objective.
+See proof in the Appendix C.
 In the distribution matching objective above, we apply the non-saturating GAN objective to minimize the adversarial divergence of marginal distributions.
@@
 where is the discriminator network.
-We find that the self-cooperative approach is connected with Consistency Training [49].
+We find that the self-cooperative approach is connected with Consistency Training (Song et al., 2023).
 However, Consistency Training considers the as an approximated ODE solution of to perform a point-to-point match.
 In contrast, our proposed objective matches and in marginal distribution level, which avoids the approximated error of ODE.
-To further ensure the mode cover of the proposed model, we can add consistency loss to our objective as regularization, which constructs our final training objective:
+To further ensure the mode cover of the proposed model, we can add consistency loss to our objective as regularization, which constructs the following loss:
+where and are pre-defined hyper-parameters.
+4 Try It On CIFAR-10 Before Scaling Up For Saving Money!
+In this section, we evaluate the performance of the proposed YOSO on CIFAR-10 (Yu et al., 2015) to verify its effectiveness under both training from scratch and fine-tuning settings.
+4.1 training strategies
+Before starting training, we introduce some effective training strategies in following for taming YOSO.
+Decoupled scheduler.
+We find that the optimal scheduler for performing consistency loss and adversarial loss are not identical.
+This is due to the cooperative adversarial loss not involving the approximated error regarding the timestep skips, enabling a substantial skip to maximize its efficacy.
+In contrast, the consistency loss is susceptible to approximated error from timestep skips, necessitating a more conservative skip to preserve its effectiveness.
+Hence, to better unleash the capability of each loss, we propose using decoupled schedulers to construct our final training objective:
+where , , and we let and in experiments.
+Model | FID | NFE |
+Distillation or Fine-tuning | ||
+YOSO | 1.81 | 1 |
+Tract (Berthelot et al., 2023) | 3.78 | 1 |
+Diff-Instruct (Luo et al., 2023c) | 4.53 | 1 |
+DMD (Yin et al., 2023) | 2.66 | 1 |
+CTM (Kim et al., 2024) | 1.98 | 1 |
+SiD (Zhou et al., 2024) | 1.92 | 1 |
+Training from scratch | ||
+YOSO | 2.26 | 1 |
+DDGANs (Xiao et al., 2022) | 3.75 | 4 |
+CT (Song et al., 2023) | 8.70 | 1 |
+iCT (Song & Dhariwal, 2024) | 2.83 | 1 |
+DDPM (Ho et al., 2020) | 3.21 | 1000 |
+EDM (Karras et al., 2022) | 2.04 | 36 |
+DDIM (Song et al., 2020) | 4.67 | 50 |
+StyleGAN2 (Karras et al., 2019) | 2.92 | 1 |
+Annealing strategy.
+Since our target is to obtain a powerful one-step denoised generation model, however, as the KL loss and consistency loss will enforce point-level matching, this may hurt performance in cases where the model capacity is insufficient.
+Therefore, we suggest reducing the weight of these two losses to zero as training progresses.
+Specifically, we let for decreasing the weight to zero with times, where is the current iterations and is the total training iterations.
+4.2 Empirical Evaluation
+Experiment Setting.
+We use the EDM (Karras et al., 2022) architecture for both UNet and discriminator.
+For the evaluation metric, we choose the Fréchet inception distance (FID) (Heusel et al., 2017).
+We train the proposed model on the CIFAR-10 dataset (Yu et al., 2015).
+We only employ MSE to compute image distance.
+Results.
+The quantitative results are shown in Tab.
+1.
+We observe that our method provides state-of-the-art performance with only a one-step manner on both training from scratch and fine-tuning settings, outperforming previous existing accelerated DMs, and GANs.
+In particular, our YOSO in fine-tuning EDM (Karras et al., 2022) even delivers better performance compared to EDM itself with ODE sampling.
+This is because our YOSO performs fine-tuning instead of distillation, the optimal solution for our YOSO is data distribution instead of teacher distribution.
 5 Towards One-Step Text-to-Image Synthesis
@@
 Latent Perceptual Loss.
-Prior research [15, 14, 49] has confirmed the effectiveness of perceptual loss in various domains.
-Notably, recent studies [26, 49] has found that the LPIPS loss [65] is crucial for obtaining few-step DMs with high sample quality.
+Prior research (Hou et al., 2017; Hoshen et al., 2019; Song et al., 2023) has confirmed the effectiveness of perceptual loss in various domains.
+Notably, recent studies (Liu et al., 2023; Song et al., 2023) have found that the LPIPS loss (Zhang et al., 2018) is crucial for obtaining few-step DMs with high sample quality.
 However, a notable drawback is that the LPIPS loss is computed in the data space, which is expensive.
-In contrast, the popular SD operate in the latent space to reduce computational demands.
+In contrast, the popular SD operates in the latent space to reduce computational demands.
 Hence, using LPIPS loss in training latent DMs is pretty expensive, as it requires not only the computation of LPIPS loss in the data space but also an additional decoding operation.
-Realizing that the pre-trained SD can serve as an effective feature extractor [58], we suggest using pre-trained SD to perform latent perceptual loss.
+Realizing that the pre-trained SD can serve as an effective feature extractor (Xu et al., 2023a), we suggest using pre-trained SD to perform latent perceptual loss.
 However, SD is a UNet, whose final layer predicts epsilon with dimensions identical to the data.
@@
 This challenge is more obvious during the initial stage of training.
-To address this issue, previous pure GANs [18] for text-to-image propose complex learning objectives and require expensive costs for training.
+To address this issue, previous pure GANs (Kang et al., 2023) for text-to-image propose complex learning objectives and require expensive costs for training.
 The learning of GANs has been shown to benefit from using a pre-trained network as the discriminator.
@@
 The advantages of the proposed strategy are twofold: 1) we use the informative pre-trained network as initialization; 2) the discriminator is defined over latent space, which is computationally efficient.
-Unlike previous work [44] that defines a discriminator over data space which required decode latent and backward from decoder, yielding expensive computational cost.
+Unlike previous work (Sauer et al., 2023c) that defines a discriminator over data space which required decode latent and backward from decoder, yielding expensive computational cost.
 By applying the Latent Discriminator, we observe a stable training process with fast convergence.
-Note that, expensive R1 regularization is not needed in our approach.
-Decoupled Scheduler.
-We find that the optimal scheduler for performing consistency loss and adversarial loss are not identical.
-Hence, to better unleash the capability of pre-trained DMs, we propose using decoupled schedulers to construct the following loss in fine-tuning pre-trained DMs:
-where , , and we let and in experiments.
 5.2 Fixing the Noise Scheduler
@@
 For example, the noise scheduler used by the SD makes the corrupted samples at the final timestep as: , where the terminal Signal-to-noise ratio (SNR) is , effectively creates a gap between training and inference.
-Previous work [23] has only observed that this makes DMs unable to produce pure black or white images.
+Previous work (Lin et al., 2024a) has only observed that this makes DMs unable to produce pure black or white images.
 However, we find that this issue yields serious problems in one-step generation.
 As shown in Fig.
-2, there are notable artifacts in one-step generation if we directly sample noise from standard Gaussian distribution.
+2, there are notable artifacts in one-step generation if we directly sample noise from standard Gaussian.
 The reason may be that in multi-step generation, the gap can be gradually fixed in sampling, while one-step generation reflects the gap more directly.
@@
 Informative Prior Initialization (IPI).
-The non-zero terminal SNR issue is similar to prior hole issues in VAEs [22, 3, 21].
+The non-zero terminal SNR issue is similar to prior hole issues in VAEs (Klushyn et al., 2019; Bauer & Mnih, 2019; Kingma et al., 2016).
 Therefore, we can use informative prior instead of non-informative prior to effectively address this issue.
@@
 As shown in Fig.
-3, we fine-tune PixArt- [6] whose terminal SNR is 4e-5 with introduced technique and -prediction, failing in one-step generation.
-Following [23], we suggest switching to v-prediction [41] and zero terminal SNR.
+3, we fine-tune PixArt- (Chen et al., 2024) whose terminal SNR is 4e-5 with introduced technique and -prediction, failing in one-step generation.
+Following (Lin et al., 2024a), we suggest switching to v-prediction (Salimans & Ho, 2022) and zero terminal SNR.
 However, we found that direct transition convergences slowly (see Appendix F).
@@
 We note that the zero terminal SNR scheduler has lower SNR than the original schedule at each timestep, formulating an effective distillation objective.
-We note that this adapt-stage converges rapidly, typically requiring only 1k iterations to initialize training YOSO.
+We note that this adapt stage converges rapidly, typically requiring only 1k iterations to initialize training YOSO.
 This enables a quick adaption to v-prediction and zero terminal SNR from pre-trained -prediction DMs, thereby comprehensively mitigating the non-zero terminal SNR issue in the noise scheduler.
@@
 In this section, we evaluate the performance of the proposed YOSO.
+Sec.
+6.1 examines YOSO in the context of text-to-image generation by fine-tuning pre-trained PixArt- (Chen et al., 2024) and Stable Diffusion (Rombach et al., 2022).
 In Sec.
-6.1, we evaluate YOSO on unconditional image generation by training from scratch, and we conduct ablation studies in Sec.
-6.2 to highlight the effectiveness of our algorithm.
-Sec.
-6.3 examines YOSO in the context of text-to-image generation by fine-tuning pre-trained PixArt- [6] and Stable Diffusion [40].
-In Sec.
-6.4, we show that the YOSO can be used for several downstream applications.
-Model | FID | NFE | Time (s) |
-YOSO | 3.82 | 1 | 0.05 |
-Ablation: | |||
-YOSO w/o consistency loss | 4.41 | 1 | 0.05 |
-YOSO w/o LPIPS loss | 4.63 | 1 | 0.05 |
-YOSO w/ | 10.85 | 1 | 0.05 |
-UFOGen [60] | 45.15 | 1 | 0.05 |
-DDGANs [54] | 3.75 | 4 | 0.21 |
-CT [49] | 8.70 | 1 | - |
-iCT [48] | 2.83 | 1 | - |
-DDPM [12] | 3.21 | 1000 | 80.5 |
-EDM [19] | 2.04 | 36 | - |
-DDIM [47] | 4.67 | 50 | 4.01 |
-StyleGAN2 w/o ADA | 8.32 | 1 | 0.04 |
-StyleGAN2 w/ADA | 2.92 | 1 | 0.04 |
-6.1 Unconditional Image Generation
-Experiment Setting.
-We use the same UNet and discriminator architecture as used in DDGANs [54].
-We set the denoising steps as four.
-For the evaluation metric, we choose the Fréchet inception distance (FID) [10].
-We train the proposed model on the CIFAR-10 dataset [63].
-Following Consistency Models [49], we employ LPIPS loss instead of MSE to compute image distance.
-Results.
-The quantitative results are shown in Tab.
-1.
-We observe that our method provides competitive performance compared with the best DMs and GANs.
-Although some variants of DMs achieve better results, they generally require much longer sample time compared to ours.
-We also find that among variants of GANs, our performance is surpassed only marginally by StyleGAN2 employing adaptive data augmentation.
-We also observe that the consistency models provide competitive and even better results compared to us.
-However, we note that the training of consistency models requires tons of computational resources and careful design of the noise schedule.
-As shown in Fig.
-4, the consistency models require much more computational resources than ours, including more parameters, larger batch size, and more training iterations.
-To ensure a fair comparison, we re-produce the consistency models by using the same batch size and training iterations as ours.
-Based on the fair setting, we find that the performance of consistency models significantly decrease, showing way worse performance compared to YOSO.
-We highlight that the low demand for computational resources enables us to scale the proposed YOSO to large-scale text-to-image datasets.
-6.2 Ablation Studies
-We provide additional insight into the effectiveness of the proposed YOSO by performing ablation studies to some potential variants on CIFAR-10 dataset.
-Effect of Consistency Loss and LPIPS Loss.
-We evaluate the effect of consistency loss and LPIPS loss in YOSO.
-We find that both consistency loss and LPIPS loss can improve the image quality, as indicated by FID (see Tab.
-1).
-However, we notice that if removing consistency loss or LPIPS loss, the generation performance would not decrease significantly.
-This indicates the effectiveness of our proposed adversarial formulation.
-Adversarial Divergence .
-The key design of our proposed YOSO is smoothing the distribution by using self-generated data instead of real data as ground truth to perform adversarial divergence.
-We evaluate the variant of using as adversarial divergence.
-As shown in Tab.
-1, the performance of the variant is significantly worse than YOSO.
-Moreover, besides sample quality, we find that the training of our proposed YOSO is much more stable than the variant.
-As shown in Fig.
-4, the discriminator loss of the variant is not stable and is much smaller than ours.
-This indicates that our approach can effectively smooth the distribution to make the distributions of fake and real samples closer to each other, making discrimination harder and hence resulting in higher discriminator loss.
-Furthermore, we observe mode collapse in the training of the variant.
-We note that a similar case also happened in SDXL-Turbo, a strong one-step text-to-image model, which also performs adversarial divergence directly against real data (See Fig.
-5).
-These results indicate the superiority of the proposed YOSO.
-Comparison to UFOGen.
-Similar to us, UFOGen does not directly perform adversarial divergence against real data either.
-However, they perform adversarial divergence over corrupted data (i.e., ).
-We argue that the formulation is less effective in constructing one-step generation model since their adversarial divergence only indirectly matches the one-step generation at clean data.
-We implement the UFOGen on CIFAR-10 with our own.
-Note that we also perform consistency loss in the UFOGen variant to ensure a fair comparison, as the consistency loss has been found useful in YOSO.
-As shown in Tab.
-1, UFOGen only obtains 45.15 FID on CIFAR-10, significantly worse than YOSO with 3.82 FID.
-This further demonstrates the superiority of our proposed YOSO.
-Through comparison with different adversarial divergence formulations, we find that our formulated adversarial divergence not only smooths the distribution to stabilize the training but also provides effective learning on one-step generation.
+6.3, we show that the YOSO can be used for several downstream applications.
+Moreover, we conduct ablation studies in Sec.
+6.2 to highlight the effectiveness of our proposed algorithm and proposed training techniques.
 6.1 Text-to-Image Generation
@@
 We initialize the GAN generator by pre-trained PixArt- which is a diffusion transformer with 0.6B parameters.
-We note that 0.6B parameters can be demonstrated as lightweight among text-to-image DMs.
 For the GAN discriminator, we construct the latent discriminator by pre-trained SD 1.5 using the proposed construction introduced in Sec.
@@
 We switch the pre-trained PixArt- to v-prediction by the proposed technique introduced in Sec.
-5.1, followed by training on the JourneyDB dataset [35] with resizing as 512 resolution.
+5.1, followed by training on the JourneyDB dataset (Pan et al., 2023) with resizing as 512 resolution.
 We apply a batch-size of 256 and a constant learning rate of 2e-5 during training.
 We observe the training convergence fast, requiring only 30k iterations and around 10 A800 days to be trained.
+We apply a batch-size of 256 and a constant learning rate of 2e-5 during training.
+We also conduct experiments on fine-tuning SD 1.5 via LoRA to show the effectiveness of the proposed YOSO, with a batch-size of 128 and a constant learning rate of 2e-5.
 Evaluation.
-For image quality, we employ Aesthetic Score (AeS) [45] to evaluate.
-We also adopt the Human Preference Score (HPS) v2.1 [53] to evaluate the image-text alignment and human preference.
-The AeS is computed by an aesthetic score predictor trained on LAION [45] datasets, without considering the image-text alignment.
+We employ Aesthetic Score (AeS) (Schuhmann et al., 2022) to evaluate image quality and adopt the Human Preference Score (HPS) v2.1 (Wu et al., 2023) to evaluate the image-text alignment and human preference.
+The AeS is computed by an aesthetic score predictor trained on LAION (Schuhmann et al., 2022) datasets, without considering the image-text alignment.
 HPS is trained to predict human preference given image-text pairs, considering the image-text alignment and human aesthetic.
-We mainly compare our model against the open-source state-of-the-art (SOTA) models, i.e., SD-Turbo [44] and SDXL-Turbo [44].
+Additionally, we include ImageReward score (Xu et al., 2024), and CLIP score (Hessel et al., 2021) to provide a more comprehensive evaluation of the model performance.
+We mainly compare our model against the open-source state-of-the-art (SOTA) models, e.g., SD-Turbo (Sauer et al., 2023c), PixArt-DMD (Chen et al., 2024; Yin et al., 2023), and Hyper-SD (a concurrent work) (Ren et al., 2024).
+Model | Backbone | HFL | Steps | HPS | AeS | IR | CS |
+Base Model | SD 2.1 | No | 25 | 27.28 | 5.66 | 0.36 | 33.46 |
+SD Turbo (Sauer et al., 2023c) | SD 2.1 | No | 1 | 27.06 | 5.31 | 0.40 | 32.21 |
+Base Model | SD 1.5 | No | 25 | 24.72 | 5.49 | 0.20 | 31.88 |
+InstaFlow (Liu et al., 2023) | SD 1.5 | No | 1 | 26.18 | 5.27 | -0.22 | 30.04 |
+Hyper-SD-LoRA (Ren et al., 2024) | SD 1.5 | Yes | 1 | 28.01 | 5.79 | 0.29 | 30.87 |
+YOSO-LoRA (Ours) | SD 1.5 | No | 1 | 28.33 | 5.97 | 0.43 | 31.33 |
+LCM-LoRA (Luo et al., 2023b) | SD 1.5 | No | 4 | 22.77 | 5.66 | -0.37 | 30.36 |
+PeRFlow (Yan et al., 2024) | SD 1.5 | No | 4 | 22.43 | 5.64 | -0.35 | 30.77 |
+TCD-LoRA (Zheng et al., 2024) | SD 1.5 | No | 4 | 22.24 | 5.45 | -0.15 | 30.62 |
+Hyper-SD-LoRA (Ren et al., 2024) | SD 1.5 | Yes | 4 | 30.24 | 5.55 | 0.53 | 31.07 |
+YOSO-LoRA (Ours) | SD 1.5 | No | 4 | 30.50 | 6.05 | 0.56 | 31.52 |
+Base Model | PixArt- | No | 25 | 31.07 | 6.12 | 1.05 | 33.81 |
+DMD (Yin et al., 2023) | PixArt- | No | 1 | 29.78 | 6.02 | 0.97 | 32.90 |
+YOSO (Ours) | PixArt- | No | 1 | 30.52 | 6.19 | 1.02 | 32.95 |
 Quantitative Results.
 The quantitative results are presented in Tab.
-3.
-To provide a direct comparison with SD-Turbo, we also fine-tune a model based on pre-trained SD 1.5 on JourneyDB with the introduced technique.
-We would like to highlight that it is better to conduct training on LAION dataset [45] which is the training set used by SD 1.5.
-However, it is not available when we conducted this research.
-Fine-tuning SD 1.5 on other datasets suffers from distribution shift, facing a risk of catastrophic forgetting.
+2.
 As shown in Tab.
-3, both YOSO-SD and YOSO-PixArt- clearly beat the SD-Turbo which is based on SD 2.1, a better SD compared to SD 1.5.
-It is also observed that YOSO-PixArt- outperforms SDXL-Turbo which is a much larger model with 2.6B parameters.
+2, YOSO clearly beat previous state-of-the-art methods across all metrics (including HPS, AeS, Image Reward score, and Clip Score).
+It is important to highlight that for the SD 1.5 backbone, we only use LoRA for fine-tuning SD 1.5, where only less than 10% parameters are tuned.
+Moroever, our method without human feedback learning (HFL) even outperforms Hyper-SD with HFL which potentially hacks the machine metric.
 Qualitative Comparison.
 To provide more comprehensive insight in understanding the performance of our YOSO, we further provide the qualitative comparison in Fig.
-5.
-The comparison shows that: YOSO clearly beats SD-Turbo in terms of both image quality and prompt alignment.
-Specifically, the samples by SD-Turbo are blurry to some extent.
-Although SDXL-Turbo with a larger model size improves the sample quality compared to SD-Turbo, we find that the samples by YOSO are sharper than SDXL-Turbo to some extent, and better in prompt alignment.
-For example, in the first row of Fig.
-5 related to a Pikachu, both SD-Turbo and SDXL-Turbo cannot follow the prompt to generate angry expressions and red eyes, while YOSO follows the prompt more precisely.
-Moreover, it is important to highlight that SDXL-Turbo seems to have a serious issue in mode collapse.
-Specifically, the third row of Fig.
-5 shows that samples by SDXL-Turbo are really close to each other.
+4.
+The comparison shows that: YOSO clearly beats existing baselines in terms of both image quality and prompt alignment, which is achieved by LoRA-finetuning.
+Specifically, the samples by SD-Turbo are blurry to some extent, and the samples by Hyper-SD are over-saturated and with notable artifacts.
+Moreover, the advantage of YOSO in text-image alignment is significant.
+For example, in the last row of Fig.
+4 related to a cow-people worker is coding, previous methods including SD 1.5 itself cannot produce samples following the prompt, while YOSO precisely follows the prompt to generate a half cow half people worker.
+Additionally, it is important to highlight that both SD-Turbo and DMD seem to have a serious issue in mode collapse.
+As shown in Fig.
+5, both SD-Turbo and DMD generate samples that are extremely close to each other.
 Overall, based on quantitative and qualitative results, we conclude that our YOSO is better in sample quality, prompt alignment, and mode cover compared with the SOTA one-step text-to-image models.
@@
 However, there is a 1024-resolution version of PixArt-, obtained by continuing training on the 512-resolution version of PixArt-.
-Motivated by the effectiveness of the LoRA combination [30], we suggest constructing a similar combination as follows:
+Motivated by the effectiveness of the LoRA combination (Luo et al., 2023b), we suggest constructing a similar combination as follows:
 where and means the model weight.
@@
 The impressive performance indicates the robust generalization ability of our proposed YOSO.
-6.3.2 Fine-tuning Stable Diffusion via LoRA
-We further conduct experiments on fine-tuning SD 1.5 via LoRA to show the effectiveness of proposed YOSO.
-As we perform LoRA fine-tuning, it is expected that it will be worse than full fine-tuning, we propose to use SDXL-Turbo to generate data for LoRA fine-tuning, which avoids the risk of distribution shift to some extent and is economically efficient.
-We call the model fine-tuned in this way as YOSO-LoRA.
-We only generate one image for each caption, avoiding the risk of mode collapse brought by SDXL-Turbo.
-We train YOSO-LoRA with only 64 batch size and 4000 iterations.
-We present the quantitative results in Tab.
-4 and qualitative results in Fig.
-6.
-It can be observed that YOSO-LoRA achieves even better results than LCM with 4 steps and is comparable to InstaFLow with full fine-tuning.
-And from Fig.
-6, YOSO-LoRA has a clearly better sample quality than LCM with 2 steps and InstaFlow.
-These results indicate that our algorithm enables one-step text-to-image synthesis with reasonable quality by LoRA fine-tuning.
+CM | Lper | IPI | Naive GAN | Coop GAN | DS | AS | HPS | AeS |
+✓ | 15.22 | 5.23 | ||||||
+✓ | ✓ | 20.08 | 5.45 | |||||
+✓ | ✓ | ✓ | 20.63 | 5.57 | ||||
+✓ | ✓ | ✓ | ✓ | 24.88 | 5.69 | |||
+✓ | ✓ | ✓ | ✓ | 26.42 | 5.85 | |||
+✓ | ✓ | ✓ | ✓ | ✓ | 27.10 | 5.94 | ||
+✓ | ✓ | ✓ | ✓ | ✓ | ✓ | 28.33 | 5.97 |
+6.2 Ablation Studies
+We provide additional insight into the effectiveness of the proposed YOSO by performing ablation studies on some potential variants on CIFAR-10 and on text-to-image generation.
+We call YOSO without decoupled scheduler and annealing strategy as YOSO-base.
+See more details in Appendix.
+Effect of Consistency Loss.
+For the effect of consistency loss in YOSO, our results suggest that it can improve the image quality, as indicated by the FID scores in Sec.
+6.1.1.
+Notably, the removal of the consistency loss does not lead to a substantial decline in generation performance, which indicates the effectiveness of our proposed adversarial formulation.
+Effect of Latent Perceptual Loss.
+To investigate the effect of latent perceptual loss proposed in Sec.
+5.1, we conduct experiments on text-to-image tasks by using it to compute the consistency loss.
+As shown in Sec.
+6.1.1, the proposed latent perceptual loss clearly improves the HPS from 15.22 to 20.08 and AeS from 5.23 to 5.45, highlighting its effectiveness.
+Effect of Decoupled Scheduler.
+We evaluate the impact of the proposed decoupled scheduler by adding it to YOSO-base.
+As evidenced by Sec.
+6.1.1 and Sec.
+6.1.1, the incorporation of the decoupled scheduler significantly enhances sample quality, as reflected by improvements in FID, HPS, and AeS metrics.
+Notably, the FID score on CIFAR-10 decreases from 3.82 to 3.30 and the HPS in text-to-image tasks is improved from 26.42 to 27.10, demonstrating the superior efficacy of the decoupled scheduler.
+Effect of Annealing Strategy.
+We also assess the effect of the proposed annealing strategy by removing it from the final formulated YOSO.
+As illustrated in Sec.
+6.1.1 and Sec.
+6.1.1, the removal of the annealing strategy degrades FID from 3.05 to 3.30 and reduces HPS from 28.33 to 27.10.
+Such results demonstrate the effectiveness of the annealing strategy in training YOSO.
+Adversarial Divergence .
+The key design of our proposed YOSO is smoothing the adversarial divergence by using self-generated data instead of real data as ground truth to perform adversarial divergence.We evaluate the variant of using as adversarial divergence.
+As shown in Sec.
+6.1.1, the performance of the variant is significantly worse than YOSO-Base, i.e., FID degrades from 3.82 to 10.85.
+Moreover, besides sample quality, we find that the training of our proposed YOSO is much more stable than the variant.
+As shown in Fig.
+7, the discriminator loss of the variant is not stable and is much smaller than ours.
+We also compare the variant in the text-to-image task.
+The results are given in Sec.
+6.1.1, from which we can see that replacing our cooperative adversarial loss with yields serious performance deterioration, i.e., HPS decreases from 26.42 to 24.88 and AeS decreases from 5.85 to 5.69.
+This indicates that our approach can effectively smooth the adversarial divergence by narrowing the gap between fake and real distribution, which makes discrimination harder and hence increases discriminator loss.
+Furthermore, we observe mode collapse in the training of this variant.
+Note that a similar case also occurs in SD-Turbo, a strong one-step text-to-image model, which performs adversarial divergence directly against real data (See Fig.
+4).
+Overall, these results indicate the superiority of the proposed YOSO.
+Comparison to UFOGen.
+Similar to us, UFOGen does not directly perform adversarial divergence against real data either.
+Instead, they employ adversarial divergence over corrupted data (i.e., ).
+However, we argue that such a formulation is less effective in constructing one-step generation model as the adversarial divergence it employs only indirectly aligns with one-step generation at the level of clean data.
+We implement UFOGen on CIFAR-10 on our own.
+Note that we also perform consistency loss in the UFOGen variant to ensure a fair comparison, as the consistency loss has been found useful in YOSO.
+As shown in Sec.
+6.1.1, UFOGen only obtains 45.15 FID on CIFAR-10, significantly worse than YOSO with 3.82 FID.
+This further underscores the superiority of our proposed YOSO.
+Through comparison with different adversarial divergence formulations, we find that our approach not only stabilizes training by smoothing the adversarial divergence but also facilitates effective learning for one-step generation.
 6.3 Application
 One promising and attractive property of DMs is that they can be used in multiple downstream tasks.
-In this section, we show the capability of our proposed YOSO in various downstream tasks, keeping the unique advantages of DMs.
-Since the relevant applications are mostly developed for SD, we mainly show the results obtained by YOSO-LoRA.
-Image-to-image. As shown in Fig. 6, our YOSO-LoRA is capable of performing high-quality image-to-image editing [32] in one step.
-Compatibility with ControlNet and Different Base Models.
-As shown in Fig.
-6, our YOSO-LoRA is compatible with ControlNet [64], following the condition well.
-Our YOSO-LoRA is also compatible with different base models (e.g., dreamshaper and realistic vision) fine-tuned from SD 1.5, preserving their style well.
+In this section, we show the capability of our proposed YOSO in various downstream tasks, keeping the unique advantages of DMs: 1) Image-to-image Editing: As shown in Fig.
+6, our YOSO-LoRA is capable of performing high-quality image-to-image editing (Meng et al., 2022) in one step; 2) Compatibility with ControlNet and Different Base Models: As shown in Fig.
+6, our YOSO-LoRA is compatible with ControlNet (Zhang et al., 2023), following the condition well.
+Our YOSO-LoRA is also compatible with different base models (e.g., Dreamshaper and Toonyou) fine-tuned from SD 1.5, preserving their style well.
 We find that when applying YOSO-LoRA to new base models, it fails to produce reasonable samples in one step.
 This may be due to the capability of LoRA and the distribution shift of training data.
-Model | NFE | HPS | AeS |
-SD 2.1-Turbo [44] | 1 | 27.06 | 5.83 |
-Realistic-vision | 50 | 29.76 | 6.35 |
-w/ YOSO-LoRA | 4 | 29.50 | 6.21 |
-w/ YOSO-LoRA-v2 | 1 | 27.13 | 6.12 |
-w/ YOSO-LoRA-v2 | 2 | 29.41 | 6.33 |
-Dreamshaper | 1 | 26.64 | 6.08 |
-w/ YOSO-LoRA-v2 |
-Enabling Data Flywheel.
-YOSO-LoRA trained on the data generated by SDXL-Turbo, fails to generate reasonable samples in one step when compatible with new base models.
-In light of the fact that combining YOSO-LoRA with new base models can generate high-quality samples in 4 steps with minimal performance loss as well as preserved style, we propose using YOSO-generated data for training YOSO-LoRA based on existing community models (e.g., Realistic-vision).
-We call the trained model YOSO-LoRA-v2.
-As shown in Tab.
-5, YOSO-LoRA-v2 can even outperform the full fine-tuned SD-2.1-Turbo in one step.
-This shows that the YOSO can be trained on slightly shifted data first, and then applied to generate high-quality un-shifted data to train a better YOSO.
-Moreover, the YOSO-LoRA-v2 can be compatible with some unseen base models (e.g., Dreamshaper) for one-step generation (see Figs.
-6 and 5).
-This shows that combining YOSO-LoRA with community models can generate unshifted training data, enabling the YOSO-LoRA-v2 to focus on learning the acceleration with better compatibility.
+7 Related Works
+Few-Step Text-to-Image Generation.
+LCM (Luo et al., 2023a) adapts consistency distillation (Song et al., 2023) for stable diffusion (SD) and enables image generation in 4 steps with acceptable quality.
+InstaFlow (Liu et al., 2023) adapts rectified flows (Liu et al., 2022; Liu, 2022) for SD, facilitating the generation of text-to-image in a mere single step.
+Despite this, the fidelity of the generated images is still poor.
+Some works (Yin et al., 2023) adapts variation score distillation (Wang et al., 2023) and diff-instruct (Luo et al., 2023c) for SD, which requires training an additional DMs for the generator distribution, akin to the discriminator in GANs.
+We found models trained in this way seem to have serious mode collapse issues (see Fig.
+5).
+Recently, combining DMs and GANs for one-step text-to-image generation has been explored.
+UFOGen (Xu et al., 2023c) extends DDGANs (Xiao et al., 2022) and SSIDMs (Xu et al., 2023b) to SD by modifying the computation of reconstruction loss from corrupted samples to clean samples.
+However, it still performs adversarial matching using corrupted samples.
+ADD (Sauer et al., 2023c) introduces a one-step text-to-image generation based on SD.
+It follows earlier research (Sauer et al., 2023a) by employing a pre-trained image encoder, DINOv2 (Oquab et al., 2024), as the backbone of the discriminator to accelerate the training.
+However, the discriminator design moves the training from latent space to pixel space, which substantially heightens computational demands.
+Moreover, they directly perform adversarial matches at clean real data, increasing the challenge of training.
+This requires the need for a better but expensive discriminator design and expensive R1 regularization to stabilize the training.
+SDXL-Lightning (Lin et al., 2024b) combines progressive distillation and adversarial training in a progressive framework.
+Recently, a concurrent work Hyper-SD (Ren et al., 2024) combines consistency distillation and adversarial training in a progressive framework, while our work can be trained end-to-end.
+However, their work performs adversarial point-level matching via consistency loss, and they still perform adversarial training via injecting noise.
+This results in less effective one-step learning, thus they rely on human feedback learning for achieving better performance.
+In contrast, we perform the adversarial training at the distribution level by replacing real data with self-generated data to smooth the adversarial divergence, which not only can stabilize the training without injecting noise but also form effective one-step generation learning.
+Moreover, compared with above mentioned diffusion-GAN hybrid models, our approaches can be trained from scratch to perform one-step generation, which is not demonstrated by them.
+Additionally, we extend our method not only to SD but also to PixArt- (Chen et al., 2024) which is based on diffusion transformer (Peebles & Xie, 2022).
+This demonstrates the wide application of our proposed YOSO.
 8 Conclusion
@@
 We plan to explore these unresolved issues in the generation field in our future work, in order to enhance the model’s capabilities in text generation, fairness, detail control, and quantitative control.
-Appendix C Safeguards
-The YOSO-SD and YOSO-PixArt- are trained on the JourneyDB dataset [35], which has undergone rigorous human and machine-based filtering to ensure that there are no harmful or violent images in the dataset.
-YOSO-LoRA and YOSO-LoRA-V2, on the other hand, were trained using data generated by SDXL-turbo and YOSO-LoRA, respectively.
-To reduce the risk associated with the models, we utilized the built-in safety checker from diffusers [51] to filter out harmful data.
+Appendix C Proof
+Proof: The cooperative adversaril loss is defined as: . Since the divergence is non-negative, and the is defined as , there exist a global optimal solution such that = 0. This completes the proof.
 Appendix D Training Algorithm
@@
 Text-to-image Diffusion Models.
-Since Diffusion models have shown a stable training process and are well-suited for scaling up generative models, numerous works have been proposed to extend DMs for text-to-image generation [39, 1, 40, 61].
-Latent DMs [40, 37] are widely adopted for high-resolution image and video generation due to their computational efficiency.
+Since Diffusion models have shown a stable training process and are well-suited for scaling up generative models, numerous works have been proposed to extend DMs for text-to-image generation (Ramesh et al., 2022; Balaji et al., 2022; Rombach et al., 2022; Xue et al., 2023).
+Latent DMs (Rombach et al., 2022; Podell et al., 2023) are widely adopted for high-resolution image and video generation due to their computational efficiency.
 Appendix F The Effect of Quick-Adaption
 We present a qualitative comparison between YOSO w/o quick adaption and YOSO w/ quick adaption in Fig.
-7.
+8.
 Note that the YOSO and variants here are trained with only 5k iterations.
@@
 As shown in Fig.
-7, YOSO w/ quick adaption demonstrates significantly better image quality compared to YOSO w/o quick adaption.
+8, YOSO w/ quick adaption demonstrates significantly better image quality compared to YOSO w/o quick adaption.
 The generated images of YOSO w/o quick adaption exhibit over-smoothing and severe artifacts.
@@
 and = 0.999.
-We adopt a constant learning rate of 2e-4 for both the discriminator and the generator.
+We adopt a constant learning rate of 2e-4 for both the discriminator and the generator in training from scratch.
+We adopt a constant learning rate of 2e-5 for both the discriminator and the generator in fine-tuning.
 We apply EMA with a coefficient of 0.9999 for the generator.
@@
 We let the and .
+For full fine-tuning, we train YOSO on the JourneyDB dataset (Pan et al., 2023), by resizing to 512 resolution.
+And we only use the square image.
+For LoRA fine-tuning, we use an internally collected dataset and the caption of the JourneyDB dataset (Pan et al., 2023) to generate data for training, we only generate one image for one caption.
+For the evaluation, we evaluate the HPS score on its benchmark, and we evaluate other metrics based on COCO-5k (Lin et al., 2014) datasets.
+G.3 Ablation Study
+For the ablation study on CIFAR-10, we use the same UNet and discriminator architecture as used in DDGANs (Xiao et al., 2022).
+For the generator, we use the Adam optimizer with = 0.9 and = 0.999; for the discriminator, we use the Adam optimizer with = 0.
+and = 0.999.
+We adopt a constant learning rate of 2e-4 for both the discriminator and the generator.
+We apply EMA with a coefficient of 0.9999 for the generator.
+We let the and .